OpenStudy (anonymous):

Evaluate without using a calculator. Answer in radians and degrees. Thank you! (:

OpenStudy (anonymous):

\[\sin^{-1} (-1)\]

OpenStudy (anonymous):

my answer was 270 degrees, (3 pi/2) however i'm terrible in math so can someone please check my answer? xD thank you.

OpenStudy (anonymous):

\[\sin ^{-1}(-1)= \Pi or 180 degrees\]

OpenStudy (anonymous):

no it is not right unfortunately because your answer needs to be in the interval \([-\pi,\pi]\)

OpenStudy (phi):

Are you guys sure?? I always thought sin(-90)= -sin(90)= -1 and -90 is 270

OpenStudy (anonymous):

mistake on my part, should have been \[\sin(-\frac{\pi}{2})=-1\] and therefore \[\sin^{-1}(-1)=-\frac{\pi}{2}\]

OpenStudy (anonymous):

@phi i was wrong it should be -90, or \(-\frac{\pi}{2}\) not \(-\pi\)

OpenStudy (phi):

@satellite73 - I was getting worried when multiple people were saying pi. btw, the answer could be the interval 0 to 2pi or -pi to +pi , depending on who's asking the question. Personally, 270 and 3pi/2 would be ok with me.